# HIRO: Hierarchical Information Retrieval Optimization

### v1

@@ Abstract
Large Language Models (LLMs) excel in natural language tasks but face limitations due to static training datasets, resulting in outdated or contextually shallow responses. Retrieval-Augmented Generation (RAG) addresses this by integrating real-time external knowledge, enhancing model accuracy and credibility, especially for knowledge-intensive tasks. However, RAG-enhanced LLMs struggle with long contexts, causing them to ”choke” on information overload, compromising response quality. Recent RAG applications use hierarchical data structures for storing documents, organized at various levels of summarization and information density. In this context, we introduce HIRO (Hierarchical Information Retrieval Optimization), a novel querying approach for RAG applications using hierarchical structures for storing documents. HIRO employs DFS-based recursive similarity score calculation and branch pruning to minimize the context returned to the LLM without informational loss. HIRO outperforms existing querying mechanisms on the NarrativeQA dataset by an absolute performance gain of 10.85%.
Figure 1: Architectural Overview of a Retrieval-Augmented Language Model (RALM). A query is inputted, processed through an embedding model to generate a vector representation, which is then matched in a vector database to retrieve relevant contexts. These contexts are fed, along with the original query, into the LLM, resulting in an informed response.
Introduction
@@
The advent of Large Language Models (LLMs) has brought about a significant transformation in artificial intelligence and natural language processing, enabling significant advancements in various tasks such as text generation, translation, and question-answering (Bubeck et al. 2023). With growing size and complexity, these models have evolved into highly effective standalone knowledge repositories, embedding vast amounts of factual information within their parameters (Petroni et al. 2019; Talmor et al. 2020). However, LLMs are inherently constrained by the static nature of their training datasets (Minaee, Mani, and Aggarwal 2024), lacking domain-specific knowledge and the adaptability to continuously evolving real-world information where the breadth and depth of knowledge required far exceed the capacity of any static training corpus. Fine-tuning, though used to tailor models, falls short in effectively integrating evolving data, especially when confronted with the need to continuously update and incorporate domain-specific or real-time information (Lewis et al. 2020). This challenge underscores a critical limitation in LLM design.
@@
In response to these limitations, the field has witnessed the emergence of Retrieval-Augmented Generation (RAG) models (Lewis et al. 2020), leading to the development of Retrieval-Augmented Language Models (RALMs). These paradigms enhance LLMs by integrating dynamic external knowledge through retrieval mechanisms. Inspired by the success of open-domain question answering systems, these approaches utilize indexed text databases, enhancing model responses by presenting retrieved information alongside input queries. RALMs not only boost the models’ ability to provide current, domain-specific knowledge but also introduce interpretability and source traceability, contrasting with the opaque parametric knowledge of traditional LLMs (Ram et al. 2023; Akyürek et al. 2022). This represents a significant stride towards overcoming the inherent limitations of static knowledge systems.
@@
Amidst the evolving landscape of information retrieval for natural language processing, novel approaches to information retrieval, such as the use of hierarchical data structures for organizing documents (Kanagavalli and Maheeja 2016; Wang et al. 2024; He et al. 2024; Sarthi et al. 2024), represents a significant step forward. Techniques like RAPTOR: Recursive Abstractive Processing for Tree-Organized Retrieval (Sarthi et al. 2024), which clusters and summarizes texts into a tree structure, showcases the potential for such methodologies to overcome traditional limitations by offering both granular and high-level insights. Yet, the adoption of hierarchical data structures presents challenges, notably the overwhelming amount of context they can return.
These hierarchical data structures typically employ fixed strategies like returning a predefined top-$k$ most relevant nodes following Tree Traversal and Collapsed Tree Querying (Sarthi et al. 2024). This approach results in a static quantity of data being retrieved, which may not align with the requirements of the query. For queries demanding extensive information, this can result in too little data being returned, whereas for simpler inquiries, it might overwhelm Large Language Models (LLMs) with excessive context. This static definition of ’k’ does not account for the variability in information needs, potentially leading to an information choke in LLMs. Overloading LLMs with too much context can degrade their ability to generate coherent and relevant responses, particularly in handling large contexts (Liu et al. 2024; Sun et al. 2021; Levy, Jacoby, and Goldberg 2024). This limitation hampers the models’ performance and undermines the efficiency and applicability of Retrieval-Augmented Language Models (RALMs) in practical scenarios.
@@
To address this challenge, our research proposes a novel querying technique designed specifically for optimizing the use of hierarchical data structures within retrieval-augmented generation systems. By applying depth-first search (DFS)-based recursive similarity thresholding and branch pruning, our approach aims to selectively minimize the context delivered to LLMs. This strategy intends to mitigate the risk of information overload, ensuring that LLMs receive a distilled yet comprehensive information set, thereby enabling more accurate and efficient response generation.
Related Work
Retirieval Augmented Generation (RAG)
@@
Retrieval-Augmented Generation (RAG) (Lewis et al. 2020), enhances Large Language Models (LLMs) by integrating real-world data from external databases, thus birthing Retrieval Augmented Language Models (RALMs). This approach ensures generated content is both contextually relevant and factually accurate, addressing the issue of LLMs producing plausible but incorrect ”hallucinated” information (Shuster et al. 2021). By enhancing the accuracy and reliability of AI-generated text, RAG heralds a new era of trustworthiness and clarity in machine-generated communications.
RAG’s widespread use in domains like information retrieval and question-answering underscores its utility. Yet, its scalability and efficiency issues, especially in handling large datasets and complex structures, highlight the crucial need for ongoing optimization research to extend its applications and enhance performance.
Adoption of Hierarchical Data Structures for Documents
@@
The adoption of hierarchical structures in RAG models, exemplified by RAPTOR (Sarthi et al. 2024), represents a pivotal shift towards more organized and efficient information storage and retrieval. RAPTOR’s approach, through recursive summarization to create a hierarchical tree, facilitates nuanced access to information at various levels of abstraction. This innovation not only improves coherence and information density tailored to specific tasks but also encourages exploration into other hierarchical frameworks like graphs for enhanced document interrelation preservation. Such advancements signal a new phase in computational linguistics, focusing on sophisticated, hierarchical data use for better knowledge representation and retrieval.
Figure 2: Tree Traversal Querying Mechanism. The top-$k$ root nodes are selected based on their similarity to the query. In the next layer, the children of these nodes are considered, and the top-$k$ nodes are again selected based on similarity. This process continues until reaching the leaf nodes.
Figure 3: Collapsed Tree Querying Mechanism. The tree is flattened into a single layer, and the top-$k$ nodes are directly selected based on their similarity to the query embedding. The returned context consists of the most relevant nodes from this single-layer selection.
Figure 4: HIRO Querying Mechanism. Our method’s recursive process, which employs Selection and Delta Thresholds to filter and prune document graphs for optimized information retrieval. It highlights how relevant contexts are dynamically identified and retrieved based on a balance of specificity and information breadth, tailored to each unique query.
Existing Querying Mechanisms
Existing Querying Mechanisms in Retrieval-Augmented Generation (RAG) models have undergone significant transformations to better address the dynamic and complex nature of queries. These models initially employed static retrieval methods, fetching a set number of documents without considering the query’s complexity. However, advancements like the Tree Traversal and Collapsed Tree Approach have introduced more adaptive methodologies (Sarthi et al. 2024).
The Tree Traversal approach progressively selects the top-$k$ most relevant nodes from each layer of the hierarchical data based on similarity. This method allows for adjusting the depth and number of nodes to control the information’s specificity and breadth.
Conversely, the Collapsed Tree approach simplifies retrieval by flattening the hierarchy, enabling simultaneous comparison of all nodes.
@@
Despite these innovations, a limitation arises from the static quantity of data retrieved, which may not always align with the query’s needs, leading to potential information overload for LLMs. This static $k$ value, irrespective of query complexity, can hamper LLMs’ performance by either providing insufficient data for complex queries or overwhelming them with excessive context for simpler ones.
Moreover, these methods tend to return both parent (summary of child nodes) and child nodes in hierarchical document structures due to matching similarity levels, this may not be efficient. This inefficiency arises from the duplicated information caused by overlapping content between the parent and child nodes. Consequently, optimizing the retrieval process to prevent redundant data retrieval becomes vital for improving the efficiency and effectiveness of LLMs in handling queries within hierarchical document structures.
Detailed pseudocode for the querying mechanisms discussed, including the Tree Traversal and Collapsed Tree approaches, is available in Appendix A1 for further reference.
Recursive Similarity Check and Branch Pruning for Querying
@@
Our approach introduces Recursive Similarity Thresholding and Branch Pruning for dynamic query response optimization. Leveraging the Similarity Threshold ($S$) and Delta Threshold ($\Delta$) hyperparameters, it tailors retrieval processes to query intricacies. $S$ filters document graphs based on similarity scores, while $\Delta$ refines relevance by pruning branches. This adaptable approach enables fine-tuning of hyperparameters at document or corpus levels, catering to specific informational needs and contexts.
Model Average Context Length
RAPTOR with Collapsed Tree Querying 2253.34586466

### v2

@@ Abstract.
Retrieval-Augmented Generation (RAG) has revolutionized natural language processing by dynamically integrating external knowledge into Large Language Models (LLMs), addressing their limitation of static training datasets. Recent implementations of RAG leverage hierarchical data structures, which organize documents at various levels of summarization and information density. This complexity, however, can cause LLMs to ”choke” on information overload, necessitating more sophisticated querying mechanisms. In this context, we introduce Hierarchical Information Retrieval Optimization (HIRO), a novel querying approach that employs a Depth-First Search (DFS)-based recursive similarity score calculation and branch pruning. This method uniquely minimizes the context delivered to the LLM without informational loss, effectively managing the challenge of excessive data. HIRO’s refined approach is validated by a 10.85% improvement in performance on the NarrativeQA dataset.
Hierarchical Information Retrieval, Retrieval-Augmented Generation (RAG), Large Language Models (LLMs)
copyright: noneccs: Information systems Information retrievalccs: Computing methodologies Search methodologiesccs: Computing methodologies Natural language processingccs: Computing methodologies Knowledge representation and reasoning
1. Introduction
@@
The advent of Large Language Models (LLMs) has brought about significant transformation in artificial intelligence and natural language processing, enabling advancements in tasks such as text generation, translation, and question-answering (Bubeck et al., 2023). As these models increase in size and complexity, they have evolved into highly effective standalone knowledge repositories, embedding vast amounts of factual information within their parameters (Petroni et al., 2019; Talmor et al., 2020). However, LLMs are inherently constrained by the static nature of their training datasets, which limits their adaptability to continuously evolving real-world information, where the breadth and depth of knowledge required far exceed any static training corpus (Minaee et al., 2024). This limitation underscores a critical gap in LLM design, where fine-tuning falls short, particularly when updates and domain-specific information are necessary (Lewis et al., 2020).
@@
In response to these challenges, the field has witnessed the emergence of Retrieval-Augmented Generation (RAG) models, evolving into what are now known as Retrieval-Augmented Language Models (RALMs). These new paradigms enhance LLMs by integrating dynamic external knowledge through sophisticated retrieval mechanisms, effectively addressing the inherent limitations of static knowledge bases (Lewis et al., 2020). Inspired by the success of open-domain question answering systems, RALMs utilize indexed text databases to enhance model responses by presenting retrieved information alongside input queries (Ram et al., 2023; Akyürek et al., 2022). This integration not only boosts the models’ ability to provide current, domain-specific knowledge but also enhances interpretability and source traceability, offering a stark contrast to the opaque parametric knowledge of traditional LLMs.
@@
Amidst this evolving landscape, novel approaches to information retrieval, such as the use of hierarchical data structures for organizing documents, represent significant advancements. Techniques like RAPTOR: Recursive Abstractive Processing for Tree-Organized Retrieval, which clusters and summarizes texts into a tree structure, demonstrate the potential to overcome traditional limitations by providing both granular and high-level insights (Kanagavalli and Maheeja, 2016; Wang et al., 2024; He et al., 2024; Sarthi et al., 2024). However, the adoption of hierarchical data structures poses challenges, particularly the overwhelming amount of context they can return, which may not always align with the query requirements (Sarthi et al., 2024). For queries demanding extensive information, this can result in insufficient data being returned, whereas for simpler inquiries, it might overwhelm LLMs with excessive context. This static definition of ’k’ in conventional querying does not account for the variability in information needs, often leading to an information ”choke” in LLMs (Liu et al., 2024; Sun et al., 2021; Levy et al., 2024).
@@
To address this, our research introduces a novel querying technique specifically designed to optimize the use of hierarchical data structures within retrieval-augmented generation systems. By implementing a depth-first search (DFS)-based recursive similarity thresholding and branch pruning, our approach aims to selectively minimize the context delivered to LLMs. This strategy intends to mitigate the risk of information overload, ensuring that LLMs receive a distilled yet comprehensive information set, thereby enabling more accurate and efficient response generation.
Figure 1. Architectural Overview of a Retrieval-Augmented Language Model (RALM). A query is inputted, processed through an embedding model to generate a vector representation, which is then matched in a vector database to retrieve relevant contexts. These contexts are fed, along with the original query, into the LLM, resulting in an informed response.
[…]
A query is inputted, processed through an embedding model to generate a vector representation, which is then matched in a vector database to retrieve relevant contexts. These contexts are fed, along with the original query, into the LLM, resulting in an informed response.
2. Related Work
@@
Retrieval-Augmented Generation (RAG) (Lewis et al., 2020), enhances Large Language Models (LLMs) by integrating real-world data from external databases, thus establishing Retrieval Augmented Language Models (RALMs). This approach ensures generated content is both contextually relevant and factually accurate, addressing the issue of LLMs producing plausible but incorrect ”hallucinated” in- formation (Shuster et al., 2021). By enhancing the accuracy and reliability of AI-generated text, RAG heralds a new era of trustworthy machine-generated communications. It’s widespread use in information retrieval and question-answering showcases its utility. However, challenges remain, particularly in scalability and efficiency when dealing with large datasets and complex structures. These issues highlight the need for ongoing research in optimization techniques, an area where Hierarchical Information Retrieval shows potential.
@@
The adoption of hierarchical structures in RAG models, exemplified by RAPTOR (Sarthi et al., 2024), represents a pivotal shift towards more organized and efficient information storage and retrieval. RAPTOR’s approach, through recursive summarisation to create a hierarchical tree, facilitates nuanced access to information at various levels of abstraction. This innovation not only improves coherence and information density tailored to specific tasks but also encourages exploration into other hierarchical frameworks like graphs for enhanced document interrelation preservation. These advancements signal a new phase in computational linguistics, focusing on sophisticated, hierarchical data use for better knowledge representation and retrieval. However, existing hierarchical models often face challenges in balancing information density and retrieval efficiency, leading to potential information overload or loss.
Dynamic Querying Mechanisms in RAG Models address the evolving nature of queries where traditional static retrieval methods, such as TF-IDF and BM25, have been complemented by more adaptive techniques. The Tree Traversal method progressively selects the top-k most relevant nodes from each layer of the hierarchical data based on similarity, allowing for adjustments in depth and number of nodes to control the information’s specificity and breadth. Conversely, the Collapsed Tree approach flattens the hierarchy to allow for simultaneous comparison of all nodes. Figure 5 and Figure 6 illustrate the workings of these algorithms, providing visual representations.
@@
Despite these innovations, a limitation arises from the static quantity of data retrieved, which may not always align with the query’s needs, leading to potential information overload for LLMs. This static k value, irrespective of query complexity, can hamper LLMs’ performance by either providing insufficient data for complex queries or overwhelming them with excessive context for simpler ones. Additionally, returning both parent and child nodes due to similar matching levels often results in inefficient, redundant information. Optimizing the retrieval process to avoid such duplication is crucial for improving LLM performance in hierarchical document structures.
@@
Our proposed approach, HIRO (Hierarchical Information Retrieval Optimization), addresses these shortcomings by employing Recursive Similarity Thresholding and Branch Pruning for dynamic query response optimization. Leveraging the Similarity Threshold ($S$) and Delta Threshold ($\Delta$) hyperparameters, HIRO customizes the retrieval process to match query complexities. The Similarity Threshold ($S$) filters document graphs based on similarity scores, while the Delta Threshold ($\Delta$) refines relevance by pruning branches. This flexible approach allows for the fine-tuning of hyperparameters at the document or corpus level to better meet specific informational needs and contexts.
Figure 2. HIRO Querying Mechanism. HIRO’s recursive process employs Selection and Delta Thresholds to filter and prune document graphs for optimized information retrieval. It highlights how relevant contexts are dynamically identified and retrieved based on a balance of specificity and information breadth, tailored to each unique query.
[…]
HIRO’s recursive process employs Selection and Delta Thresholds to filter and prune document graphs for optimized information retrieval. It highlights how relevant contexts are dynamically identified and retrieved based on a balance of specificity and information breadth, tailored to each unique query.
3. Methodology
Our approach introduces Recursive Similarity Score Calculation and Branch Pruning to optimize query responses through dynamic retrieval processes customised to the specifics of each query. This method is grounded in the utilization of two hyperparameters: the Selection Threshold ($S$) and the Delta Threshold ($\Delta$), which combine to create a dynamic retrieval process depending on the complexity of the query.
The detailed pseudocode for our HIRO approach is outlined in Algorithm 1. For a version optimized for deployment in production environments, please refer to Algorithm 3 in Appendix B, which includes both the primary HIRO algorithm and the auxiliary Algorithm 4 for evaluating child nodes.
3.1. Recursive Similarity Score Calculation
The Selection Threshold (S) serves as an initial filter, identifying document graphs for further exploration based on similarity scores between the query and parent nodes. Only graphs with parent nodes surpassing this threshold proceed, ensuring focus on potentially relevant content.
The process begins at the root layer of the hierarchical data structure, where the similarity between the query embedding $Q$ and the embeddings of all nodes $N_{i}$ present at this initial layer is computed. The similarity score $S(Q,N_{i})$ is determined using cosine similarity, which measures the cosine of the angle between two non-zero vectors, providing a value between -1 and 1:
$S(Q,N_{i})=\frac{Q\cdot N_{i}}{\|Q\|\|N_{i}\|}$
Nodes whose similarity scores exceed the predefined Selection Threshold $S$ are earmarked for further exploration. This step ensures that only nodes with a significant level of relevance to the query are considered in subsequent stages.
Earmarked $\text{Earmarked Nodes}=\{N_{i}\mid S(Q,N_{i})>S\}$
3.2. Branch Pruning
The Delta Threshold ($\Delta$) refines the search by evaluating each child node within accepted graphs. This step is crucial in eliminating less relevant data, thus enhancing the efficiency and relevance of the retrieval process.
For each earmarked parent node $N_{i}$, the similarity between the query embedding $Q$ and each child node $N_{j}$ is calculated. The child nodes’ similarity scores $S(Q,N_{j})$ are compared to their respective parent nodes’ scores. The difference in similarity, denoted as $\Delta S$, is computed as follows:
$\Delta S=S(Q,N_{j})-S(Q,N_{i})$
A child node is marked for further recursive evaluation if $\Delta S$ exceeds the Delta Threshold $\Delta$:
$\Delta S>\Delta$
If a child node does not meet this criterion, the parent node is retained for context. This selective pruning ensures that only the most promising branches are explored, enhancing the efficiency of the retrieval process.
3.3. Dynamic Retrieval Process
The dynamic retrieval process aggregates the selected nodes to form an optimized and coherent context for the language model (LLM). This approach balances the retention of pertinent information with the need to avoid overwhelming the system with excessive detail.
3.3.1. Termination Criteria
The recursive search terminates under two conditions: when the current node is a leaf node, or when no further nodes meet the threshold criteria. Specifically, the search ends if the current node does not have any children or if the similarity score improvement does not exceed the Delta Threshold.
3.3.2. Context Aggregation
Upon termination of the search, the selected nodes are aggregated to form the final context for the LLM. Nodes $N_{k}$ are included if their similarity $S(Q,N_{k})$ to the query exceeds the Selection Threshold $S$ and if each child node $N_{j}$ shows a similarity gain $\Delta S_{kj}=S(Q,N_{j})-S(Q,N_{k})$ above the Delta Threshold $\Delta$. This ensures the LLM receives a coherent and comprehensive set of relevant information.
Final $\text{Final Context}=\left\{N_{k}\mid S(Q,N_{k})>S\text{ and }\forall N_{j},\,\Delta S_{kj}>\Delta\right\}$
3.4. Hyperparameter Impact
The Selection Threshold $S$ governs the scope of data examined, controlling the breadth of information explored. In contrast, the Delta Threshold $\Delta$ governs the extent of information examined, managing the depth of the search.
The interplay between $S$ and $\Delta$ is critical:
  • Selection Threshold $S$: A higher $S$ means only highly similar nodes are considered, narrowing the initial search but potentially missing relevant sub-nodes.
  • Delta Threshold $\Delta$: A higher $\Delta$ means only significant improvements in similarity are considered, reducing the depth of search and focusing on more promising nodes.
By fine-tuning these hyperparameters, the retrieval process can be dynamically adapted to the complexity and requirements of each specific query, ensuring efficient and relevant information extraction.
Algorithm 1 HIRO Querying
  Output: Context Nodes
1:  $context\leftarrow\text{[]}$]
[…]
3:  while $nodes\_to\_evaluate$ is not empty do
[…]
5:     if $node$.parent exists then
6:        $parent\_score\leftarrow\text{similarity}(query,node.parent)$
7:     else
8:        $parent\_score\leftarrow\text{0}$
9:     end if
10:     $score\leftarrow\text{similarity}(query,node)$
11:     $delta\leftarrow score-parent\_score$
12:     if (($score$ $S$ and $node$.is_leaf()) or ($delta$ $\Delta$)) then
13:        $context$.append($node$.content)
14:     else
[…]
16:     end if
17:  end while
18:  return $context$
4. Experiments
4.1. Datasets
We assessed the performance of HIRO paired with RAPTOR using two prominent question-answering datasets: NarrativeQA and QuALITY.
NarrativeQA is an extensive English-language dataset that includes 1,567 stories, drawn from both books and movie scripts. The dataset is accompanied by 46,765 questions, which are crafted based on human-generated abstractive summaries of the stories. The stories range between 50,000 and 100,000 words (Kočiský et al., 2018). This dataset is designed to evaluate a system’s ability to focus on key narrative elements while filtering out irrelevant details. It tests systems’ ability to prune irrelevant details for key story elements, showcasing querying optimization for narrative understanding and information retrieval. Evaluation metrics include ROUGE, which measures recall of n-grams; BLEU-1 and BLEU-4, which assess the precision of n-grams in generated summaries; and METEOR, which evaluates precision and recall by considering synonyms and stemming.
The QuALITY dataset (Question Answering with Long Input Text, Yes!) is designed for evaluating multiple-choice question answering tasks on lengthy English texts. It contains articles with lengths ranging from 2,000 to 8,000 tokens, providing a diverse range of textual complexity (Pang et al., 2021). The dataset includes 6,737 questions, with a challenging subset named QuALITY-HARD, which features 3,360 particularly difficult questions. This subset is intended to assess the capability of QA models to extract and utilize relevant information from extensive text passages effectively. Performance is measured using the F1 score, which balances precision and recall, and Accuracy, which determines the proportion of correct answers. These metrics help evaluate how well different pruning methods and information retrieval strategies perform in identifying crucial details within long documents.
4.2. Controlled Baseline Comparisons
To establish baseline performance, we utilize RAPTOR with SBERT (Reimers and Gurevych, 2019) as the embedding model and GPT-3.5 Turbo (Brown et al., 2020) as the reader/language model. This configuration represents the state-of-the-art approach in retrieval-augmented generation using RAPTOR. We systematically compare the performance of RAPTOR combined with SBERT and GPT-3.5 Turbo, both with and without the implementation of HIRO querying.
Baseline Configuration: Performance is evaluated using RAPTOR with SBERT and GPT-3.5 Turbo. RAPTOR traditionally employs a method known as Collapsed Tree Querying, where it flattens the hierarchical tree structure and selects the top-$k$ nodes based on similarity scores. This conventional approach is compared against the HIRO-enhanced method, which incorporates hierarchical querying to optimize the retrieval process.
Comparison with HIRO Querying: We analyze the performance difference between the conventional Collapsed Tree Querying approach and the HIRO querying method. The primary focus is on understanding how HIRO’s dynamic traversal and adaptive thresholds improve retrieval efficiency and information relevance compared to the static approach of selecting top-$k$ nodes.
Token Length and Computational Trade-offs: We also examine the token length of the final context returned to the language model (LM) after querying. This analysis provides insights into the trade-off between computational resources and performance gains. By evaluating how the length of the context affects performance and computational load, we aim to understand the efficiency of different querying methods in managing token limits and processing requirements.
Non-linear Performance Degradation: Research indicates that the performance of large language models (LLMs) degrades non-linearly with increasing prompt token length (Levy et al., 2024). To account for this, we present normalized performance scores for both conventional and HIRO querying methods. These scores are adjusted by the logarithm of the context length (in tokens) to offer a clearer picture of efficiency relative to context size. This normalization helps in comparing the effectiveness of querying methods while considering the impact of prompt length on model performance.
Computational Time Complexity: Finally, we present the time complexity analysis for the conventional Tree Querying Methods and HIRO. This analysis highlights the computational cost associated with each approach, providing a detailed comparison of their efficiency.
4.3. Experimental Setup
4.4. Experimental Results
Model Performance on Sequence Generation Tasks: Table 1 shows the average context length in tokens returned by different querying mechanisms with the RAPTOR model on the NarrativeQA dataset. The HIRO querying mechanism strikes an optimal balance by returning shorter contexts than tree traversal but longer than collapsed tree querying, suggesting an effective trade-off between information overload and sufficiency.
Table 1. Comparison of Average Context Length (in tokens) returned by Tree Traversal, HIRO and Collapsed Tree Querying on the NarrativeQA Dataset.
Model Configuration Context Length